### v6

# Gradient-Guided Annealing for Domain Generalization

Abstract
Domain Generalization (DG) research has gained considerable traction as of late, since the ability to generalize to unseen data distributions is a requirement that eludes even state-of-the-art training algorithms. In this paper we observe that the initial iterations of model training play a key role in domain generalization effectiveness, since the loss landscape may be significantly different across the training and test distributions, contrary to the case of i.i.d. data. Conflicts between gradients of the loss components of each domain lead the optimization procedure to undesirable local minima that do not capture the domain-invariant features of the target classes. We propose alleviating domain conflicts in model optimization, by iteratively annealing the parameters of a model in the early stages of training and searching for points where gradients align between domains. By discovering a set of parameter values where gradients are updated towards the same direction for each data distribution present in the training set, the proposed Gradient-Guided Annealing (GGA) algorithm encourages models to seek out minima that exhibit improved robustness against domain shifts. The efficacy of GGA is evaluated on five widely accepted and challenging image classification domain generalization benchmarks, where its use alone is able to establish highly competitive or even state-of-the-art performance. Moreover, when combined with previously proposed domain-generalization algorithms it is able to consistently improve their effectiveness by significant margins [^1] Code available at: https://github.com/aristotelisballas/GGA.
[Uncaptioned image][Uncaptioned image]
Figure 1: (left) Decision boundaries of a 4th-degree polynomial logistic regression model with 2D input. In this example, feature $x_{1}$ is class-specific and $x_{2}$ is domain-specific, while color represents classes and shapes represent domains. The samples with solid red and green colors are included in the training data, whereas the fainted samples are part of the hidden held-out test set. As a result, domain shift is represented by a change in $x_{2}$. Although the classifier should only infer based on $x_{1}$, traditional gradient descent leads to overfitting (top-left). The proposed method, GGA (bottom-left), introduces an annealing process that depends on gradient agreement, leading to models that generalize well to new, unobserved target domains. (right) Schematics of the parameter updates of ERM (top-right) and GGA (bottom-right). Parameters updated via ERM are driven by gradient conflict, whereas GGA searches for a point where gradients align before continuing descending towards a minima.
1 Introduction
The vast majority of neural networks today are trained via stochastic gradient descent methods, such as SGD [12] or ADAM [40], where the gradient direction guides the optimization through the loss landscape, aiming to converge to a minimum. What is more, it has been empirically shown that the astounding performance and generalization capabilities of these over-parameterized models stems from the fact that loss surfaces of large neural networks have multitudinous local minima [2, 65], most of which yield similar performance upon model convergence [20, 57].
The above optimization process assumes that all available data samples are independent and identically distributed (i.i.d.). When this assumption holds, the parameter values reached during training is likely to generalize to the test distribution. However, in several real-world scenarios the i.i.d. assumption is not met, and models are evaluated on similar but distinct out-of-distribution (OOD) data resulting from domain shifts to the training distribution, leading to a domain generalization (DG) [96] problem. In this case, the training loss minimum may lead to a poor parameter configuration for the test data.
The disagreement of gradients across data from different domains during training, can provide an indication that the described problem occurs. This observation about gradient conflicts was initially made in the multi-task learning paradigm [92], where gradients of different tasks pointed to conflicting directions on the loss surface. To mitigate the effects of gradient conflicts, methods to balance the relative gradient magnitudes were proposed [18, 38], along with algorithms that remove disagreeing gradient components among tasks [68, 86]. Similarly, gradient conflicts were also addressed in the context of DG in [53], where gradients with different signs among domains were either muted or set to random values; as they are assumed to contain domain-specific information. With that being said, it has been empirically shown that even though the above issues arise, the correct selection and tuning of hyperparameters can be enough for vanilla network training to surpass even state-of-the-art algorithms [31]. This evidence leads us to believe that there exist local minima on the vanilla loss surface that lead to more robust and generalizable models. Depending on the starting parameter configuration, even a model optimized by traditional Empirical Risk Minimization can reach a solution that is locally optimal for all distinct domains. These hypotheses, along with problems poised by gradient disagreement, are demonstrated in Section 1.1 (illustrated in Fig. 1) and are further explored in Section 3.2.
Starting from these ideas, we propose an alternative strategy for updating the parameters of a neural network during the optimization process, in an attempt to “set the model up for success”. Specifically, inspired by Simulated Annealing Optimization [41], during the early stages of model training we iteratively anneal, i.e. randomly perturb the parameters of the model, and search for a set of parameter values where the gradients between all training domains agree, before minimizing the total loss across all training domains. We call this simple, yet effective, strategy Gradient-Guided Annealing or GGA. When evaluated on extensive and challenging DG benchmarks, GGA is able to boost the performance of the baseline by significant margins, even yielding state-of-the-art results without the application of additional data processing or augmentation techniques. Additionally, since GGA can be considered a general strategy for handling multi-domain datasets it can also be combined with previously proposed algorithms. Experimental results of the combined methods, demonstrate the efficacy of GGA, as it is able to enhance their generalization capacity and overall accuracy, boosting their performance over the baseline benchmarks.
Our primary contributions are as follows:
  • We present Gradient-Guided Annealing (GGA), a DG method for training neural networks such that gradients align across domains.
  • We validate the effectiveness of GGA on multiple, challenging Domain Generalization benchmarks both as a standalone algorithm and by combining it with previous state-of-the-art methods.
  • We offer further evidence on the effectiveness of the proposed method by investigating the domain gradients during training and its sensitivity to the choice of hyperparameters.
1.1 Gradient Disagreement and Domain Overfitting
To demonstrate how the domain generalization problem manifests during optimization in the source domains during training, we consider the following simple synthetic binary classification problem. For training, data is sampled from a mixture of two domains, while testing takes place on a different, previously unseen, target domain. Each sample has two features, a class-specific feature $x_{1}$ and a domain-specific feature $x_{2}$. Our goal is to learn a model on the source domains, that can discriminate between classes effectively on the unseen target domain.
Concretely, for the source domains, we draw 200 points from a 2-D Gaussian distribution with an isotropic covariance matrix for each domain and class, i.e.,
$\bm{x}_{y}^{(d)}\sim\mathcal{N}(\bm{\mu}_{y}^{(d)},\sigma\bm{I}_{2})$ (1)
Where the subscript $y$ indicates the class and the exponent $d$ the domain, while $\bm{\mu}_{1}^{(1)}=[-2.5,-2.5]$, $\bm{\mu}_{1}^{(2)}=[-2.5,2.5]$, $\bm{\mu}_{2}^{(1)}=[2.5,-2.5]$, $\bm{\mu}_{2}^{(2)}=[2.5,2.5]$, $\sigma=0.5$ and $\bm{I}_{2}$ is the $2\times 2$ identity matrix. We also draw an additional 400 samples from a held-out test domain with means at $\bm{\mu}_{1}^{(3)}=[-2.5,-7.5]$ and $\bm{\mu}_{2}^{(3)}=[2.5,-7.5]$. The drawn samples are shown in the left column of Fig. 1. In this example, the classes can be distinguished solely on $x_{1}$ (the “class-specific” feature), while domains differ in terms of feature $x_{2}$ (the “domain-specific” feature).
We train a 4th degree polynomial logistic regression model trained with Empirical Risk Minimization (ERM, [74]) and binary cross-entropy loss using the SGD optimizer. In the top-left example of Fig. 1 the initial parameter conditions were such that training converged to a local minimum of the loss that leads the model to consider both $x_{1}$ and $x_{2}$ for its decisions. In this case, the model has clearly overfit its source domains and will fail when presented with out-of-distribution data from the held-out domain. An indicator of this was the fact that gradients of the loss for samples of different domains were dissimilar during training. This is in contrast to the bottom-left GGA model in Fig. 1 which mostly relies on $x_{1}$ to discriminate between classes. This model was trained by using the proposed gradient agreement strategy and although the resulting training loss is slightly higher, the model successfully generalizes to new domains.
In the rest of the paper, we first discuss the most relevant works in the domain generalization literature (Section 2) and then present the proposed methodology (Section 3). Followingly, we present the experimental setup and results (Section 4), and finally conclude the paper with a discussion on limitations of our method and directions for future research (Section 5).
2 Related Work
Domain Generalization (DG) methods focus on learning a model from one or multiple source data sets, or domains, which can generalize to previously unseen, out-of-distribution target domains. Existing DG methods in the literature can be categorized into two major groups; single-source and multi-source. In addition to not having any knowledge about the unseen data, single-source algorithms do not leverage information regarding the presence of disinct domains in the training set. On the other hand, multi-source methods utilize domain labels and often take advantage of the statistical differences in the sample distributions. Specifically, most popular algorithms include data augmentation [95, 14] which proves beneficial for regularizing over-parameterized neural networks and improving generalization, meta-learning [48, 91, 3, 24], which exposes models to domain shifts during training, and disentangled representation learning [60, 78, 89, 5], where models most commonly include modules or architectures that focus on decomposing learned representations into domain-specific and domain-invariant parts. Additionally, domain alignment [55, 29, 83] and causal representation learning algorithms [52, 51] have also been proposed in the literature towards producing robust models that retain their generalization capabilities on unseen data. Finally, ensemble learning methods [97] have also been proposed for DG, where techniques such as weight averaging [37] lead to improved generalization [15].
Gradient operations for DG. Lately, there has been a surging interest in addressing the DG problem from a gradient-aware perspective. The most relevant works to ours, leverage gradient information to learn generalized representations from the source datasets. For example, [34] proposes a self-challenging learning scheme, by muting the feature representations associated with the highest gradients and forcing the network to learn via alternative routes. In another work, the authors of [71] propose Fish, a first-order algorithm that approximates the optimization of the gradient inner product between domains. Inspired by gradient surgery in multi-task learning [86], [53] proposes aligning the gradient vectors among source domains by retaining the individual same-sign gradients, and either setting the rest to zero or random values. Finally, there has also been great interest into researching the properties of Sharpness-Aware Minimization (SAM) [28, 98, 81, 45], as it has been hypothesized that flatter minima lead to smaller DG gaps and improved generalization.
Simulated Annealing for Deep Learning. Although explored in past literature, simulated annealing (SA) has not been explicitly proposed for DG. To avoid being “trapped” in local minima during model optimization, [13] proposes SA-GD, a simulated annealing method for gradient descent. Similarly, [63] shows that by sacrificing computation time, simulated annealing optimization can also improve the results of standard CNN architectures. In a more recent work, the authors of [67] propose SEAL and apply simulated annealing in the early layers of networks to prohibit them from learning overly specific representations and improve model generalization. Furthermore, there has also been research regarding the combination of SA with reinforcement learning (RL) algorithms, where RL is used to optimize specific hyperparameters of the SA process [21].
When compared to previously proposed methods which take into consideration gradient behaviour, GGA has some key differences. In contrast to gradient surgery algorithms [86, 53] that either mute, aggregate or set gradients to random values, and Fish [71] that approximately optimizes the inner-product between domain gradients, GGA searches for existing parameter space points where gradients of different domains have pairwise small angles. Furthermore, as GGA is applied in the early stages of training and for a limited number of training iterations, the high computational burden of traditional simulated annealing methods is avoided.
3 Methods
3.1 Preliminaries
Consider a classification problem with $K$ classes. During model training we have access to a data set composed of distinct source data distributions (or domains), $\mathcal{S}=\left\{D_{1},D_{2},\dots,D_{\lvert\mathcal{S}\rvert}\right\}$. From each domain $D_{i}$, we observe $n_{i}$ labeled data points, such that $(\bm{x}_{j}^{(i)},y_{j}^{(i)})\sim D_{i}$, for $j=1,\dotsc,n_{i}$. Similarly, the test dataset consists of $\mathcal{T}=\left\{D^{T}_{1},D^{T}_{2},\dotsc,D^{T}_{\lvert\mathcal{T}\rvert}\right\}$ unseen target data distributions, from which the model cannot retrieve any information during training. The goal is to learn a single labeling function $h(\bm{x};\bm{\theta})$, parameterized by $\bm{\theta}$, which correctly maps input observations $\bm{x}_{j}^{(i)}$ to their labels ${y_{j}^{(i)}}$ for both the seen source and the unseen target domains.
If $\mathcal{L}_{i}(\bm{\theta})=\frac{1}{n_{i}}\sum_{j=1}^{n_{i}}\ell(h({\bm{x}_{j}}^{(i)};\bm{\theta}),{y_{j}^{(i)}})$ represents the loss associated to the $i$-th source data domain in the training set, we define the overall cost function $\mathcal{L}(\bm{\theta})=\frac{1}{\lvert\mathcal{S}\rvert}\sum_{i=1}^{\lvert\mathcal{S}\rvert}\mathcal{L}_{i}(\bm{\theta})$, as the average loss over all available source domains. The function $\ell(\cdot,\cdot)$ is a classification loss, in our case cross-entropy, that measures the error between the predicted label $\hat{y}$ of an input observation and its true label $y$. The standard way of model training is Empirical Risk Minimization or ERM which uses the following objective on the training domain data
$\bm{\theta}^{*}=\arg_{\bm{\theta}}\min\frac{1}{\lvert\mathcal{S}\rvert}\sum_{i=1}^{\lvert\mathcal{S}\rvert}\mathcal{L}_{i}(\bm{\theta})+\lambda R(\bm{\theta})$ (2)
where $R(\cdot)$ is a regularization term and $\lambda$ a hyperparameter responsible for controlling the contribution of regularization to the loss, leading to a parameter vector $\bm{\theta}^{*}$. In practice, networks trained via ERM have been shown to overfit the data distributions present in the training set. Previous works (such as [71]) have observed that the directions of gradients for different domains during training play a significant role in model generalization. Given source domain losses $\mathcal{L}_{i}(\bm{\theta})$ and $\mathcal{L}_{j}(\bm{\theta})$ and their corresponding gradients $\bm{g}_{i}=\nabla_{\bm{\theta}}\mathcal{L}_{i}(\bm{\theta})$, $\bm{g}_{j}=\nabla_{\bm{\theta}}\mathcal{L}_{j}(\bm{\theta})$, gradient conflicts arise when the angle of the gradients grows e.g., above $\lvert\pi/2\rvert$, or, equivalently, their cosine similarity $\frac{\bm{g}^{T}_{i}\cdot\bm{g}_{j}}{\|\bm{g}_{i}\|\|\bm{g}_{j}\|}$ becomes negative. In the following section, we further explore the role of gradient similarity as an indicator for domain overfitting.
3.2 A Generative Model for Domain Generalization
$\mathbf{x}$$f$$\mathbf{e}$$\mathbf{z}_{y}$$\mathbf{z}_{d}$$y$$d$
Figure 2: Simplified generative model for multi-domain data.
To assist with the development and understanding of the proposed method consider the generative model of Figure 2, summarized by the following process:
$y$ $\sim Multinoulli(p_{1},\dotsc,p_{K})$
$d$ $\sim Multinoulli(q_{1},\dotsc,q_{L})$
$\bm{z}_{y}$ $\sim p_{y}$
$\bm{z}_{d}$ $\sim p_{d}$
$\bm{e}$ $\sim p_{e}$
$\bm{x}$ $=f(\bm{z}_{y},\bm{z}_{d},\bm{e})$
where $y$ is a variable corresponding to the class, $d$ is a variable corresponding to the domain (with $L=\lvert\mathcal{S}\rvert+\lvert\mathcal{T}\rvert$), $\bm{z}_{y}$ is a latent multivariate class-specific representation of class $y$, drawn from an unknown distribution $p_{y}$, and $\bm{z}_{d}$ is a latent multivariate domain-specific representation of the domain $d$, drawn from an unknown distribution $p_{d}$. Notice that these representations are disentangled, i.e., $\bm{z}_{y}$ does not depend on $d$. The observed sample $\bm{x}$, on the other hand, is derived using the function $f(\bm{z}_{y},\bm{z}_{d},\bm{e})$ that depends on both the class and domain representations, as well as independent nuisance variables, $\bm{e}$.
Without loss of generality, let’s assume that we have two source domains ($\lvert\mathcal{S}\rvert=2$) and two classes ($K=2$). Then, during model training we sample only from the source domains and use the loss $\mathcal{L}=\alpha_{1}\mathcal{L}_{1}+\alpha_{2}\mathcal{L}_{2}$,
$\mathcal{L}_{i}=-\mathbb{E}_{d=i}\left[\log h(y|\bm{x};\bm{\theta})\right]$ (3)
where $h$ is the probability estimated by our model for class $y$, $\alpha_{i}$ is the percentage of samples of domain $i$, and $i=1,2$. The expectation is over samples that have been generated using the above process when $d=i$. Each gradient update step depends on the loss gradient
$\begin{split}\nabla_{\bm{\theta}}\mathcal{L}&=\alpha_{1}\nabla_{\bm{\theta}}\mathcal{L}_{1}+\alpha_{2}\nabla_{\bm{\theta}}\mathcal{L}_{2}\\ &=-\alpha_{1}\mathbb{E}_{d=1}\left[\nabla_{\bm{\theta}}\log h(y|\bm{x};\bm{\theta})\right]\\ &-\alpha_{2}\mathbb{E}_{d=2}\left[\nabla_{\bm{\theta}}\log h(y|\bm{x};\bm{\theta})\right]\end{split}$ (4)
where we have used the linearity of the expectation to obtain the expectation of the derivatives. Notice that the gradient update step depends on $\nabla_{\bm{\theta}}\log h(y|\bm{x};\bm{\theta})=\nabla_{\bm{\theta}}\log h(y|f(\bm{z}_{y},\bm{z}_{d},\bm{e};\bm{\theta}))$.
For models that use an internal domain-invariant representation (i.e., a representation that does not depend on $\bm{z}_{d}$) it will hold that $h(y|\bm{x})=h(y|f(\bm{z}_{y},\bm{z}_{1},\bm{e}))=h(y|f(\bm{z}_{y},\bm{z}_{2},\bm{e}))$, and therefore $\mathbb{E}_{d=1}\left[u(h(y|\bm{x}))\right]=\mathbb{E}_{d=2}\left[u(h(y|\bm{x}))\right]$ for any function $u$. Thus, both the losses, $\mathcal{L}_{i}$ and their gradients, $\nabla_{\bm{\theta}}\mathcal{L}_{i}$, should be equal, in expectation, for different domains $i$. This observation provides a necessary condition for domain-invariance, which in turn provisions the main motivation for the development of our method, presented in the following section.
3.3 Gradient-Guided Annealing for DG
Algorithm 1 Implementing GGA in training
1:Pretrained DNN $h$ with parameters $\bm{\theta}_{0}$, training dataset with domains $\mathcal{S}=\{D_{i}\}_{1}^{\lvert\mathcal{S}\rvert}$. Loss gradients $\bm{g}$. Learning rate $\eta$, perturbation parameter $\rho$, optimization step to start the annealing process $A_{s}$, optimization step to end the annealing process $A_{e}$, number of annealing iterations per optimization step $n_{a}$. Total number of training iterations $n$.
2:
3:for $t\leftarrow 1$ to $n$ do
4:     Sample a mini-batch: $\mathcal{B}\leftarrow\mathcal{B}_{\mathcal{D}_{1}}+...+\mathcal{B}_{\mathcal{D}_{\lvert\mathcal{S}\rvert}}$
5:     if $A_{s}\leq t\leq A_{e}$ then
6:         #Begin Gradient-Guided Annealing:
7:         Compute the mini-batch loss at starting params:
8:              $\mathcal{L}_{B}\leftarrow\mathcal{L}(\mathcal{B};\bm{\theta}_{t})$
9:         Calculate minimum domain grad pair sim:
10:          $\text{sim}\leftarrow\mathop{\min\limits_{\begin{subarray}{c}i\leq\lvert\mathcal{S}\rvert,\,j\leq\lvert\mathcal{S}\rvert\\ i\neq j\end{subarray}}}\left(\frac{\bm{g}^{T}_{i}\cdot\bm{g}_{j}}{\|\bm{g}_{i}\|\|\bm{g}_{j}\|}\right)$
11:         $\bm{\theta}_{t}\leftarrow\bm{\theta}_{t-1},\bm{\theta}^{\prime}_{t}\leftarrow\bm{\theta}_{t-1}$
12:         for  step $a\leftarrow 1$ to $n_{a}$ do
13:              $\bm{\theta}_{a}\leftarrow\bm{\theta}_{t}+\mathcal{U}(-\rho,\rho)$
14:              Calculate minimum domain grad pair sim:
15:              $\text{sim}_{a}\leftarrow\mathop{\min\limits_{\begin{subarray}{c}i\leq\lvert\mathcal{S}\rvert,\,j\leq\lvert\mathcal{S}\rvert\\ i\neq j\end{subarray}}}\left(\frac{\bm{g}^{T}_{i}\cdot\bm{g}_{j}}{\|\bm{g}_{i}\|\|\bm{g}_{j}\|}\right)$
16:              Compute the mini-batch loss at new params:
17:              $\mathcal{L}_{a}\leftarrow\mathcal{L}(\mathcal{B};\bm{\theta}_{a})$
18:              if $\big{(}\text{sim}_{a}>\text{sim}\big{)}\wedge\big{(}\mathcal{L}_{a}-\mathcal{L}_{\mathcal{B}}<0.1\big{)}$ then
19:                  $\mathcal{L}_{\mathcal{B}}\leftarrow\mathcal{L}_{a}$, $\bm{\theta}^{\prime}_{t}\leftarrow\bm{\theta}_{a}$, $\text{sim}\leftarrow\text{sim}_{a}$
20:              end if
21:         end for
22:     end if
23:     #Update weights:
24:     Compute mini-batch loss: $\mathcal{L}_{\mathcal{B}}=\mathcal{L}(\mathcal{B};\bm{\theta}^{\prime}_{t})$
25:     $\bm{\theta}_{t+1}=\bm{\theta}^{\prime}_{t}-\eta\cdot\nabla\mathcal{L}_{\mathcal{B}}(\mathcal{B};\bm{\theta}^{\prime}_{t})$
26:     $t=t+1$
27:end for
The analysis presented in the previous section shows that we can use the agreement of gradients $\nabla\mathcal{L}_{i}$ as an indicator of domain invariance, since the less the model $h(\cdot|\bm{\theta})$ depends on the domain $d$, the more similar the expected value of the gradient, $\mathbb{E}[\nabla\mathcal{L}_{d}]$, will be (see Eq. (4)). Inspired by Simulated Annealing [41], we achieve this by adding random noise to the parameters of the model in search for a point with high domain gradient agreement, as measured by the increase of the minimum gradient similarity across any pair of domains,
[…]
In more detail, as it is common in DG literature, we start with a pre-trained model $f(\cdot,\bm{\theta}_{0})$ and perform a small number of warmup training steps. This ensures that the model approaches a region of the parameter space that achieves a low loss for the target problem. We then calculate the minimum pairwise gradient similarity among source domains and begin searching the neighborhood of the parameter space for points where both: (a) the domain gradients agree and (b) the loss are reduced [^2] In our experiments, we relax the loss constraint and accept the new parameter set when the gradient similarity has increased and training loss has been reduced within a specific range (i.e $\mathcal{L}-\mathcal{L}^{\prime}<0.1$), where $\mathcal{L}^{\prime}$ is the loss value at the updated set of parameters.. This is implemented through iterative random perturbations $\bm{\theta}^{\prime}\leftarrow\bm{\theta}+\mathcal{U}(-\rho,\rho)$, where $\mathcal{U}$ is the multivariate uniform distribution. Each new point is selected only if it simultaneously achieves higher gradient agreement and lower loss (thus improving the Pareto front). After a fixed number of iterations, the point that has achieved the highest minimum gradient agreement and the lowest loss is selected. This process is repeated after every training iteration step. After a number of optimization steps, we allow the model to be trained without parameter perturbations, following a standard SGD-based procedure, as usual. The GGA algorithm is presented in Algorithm 1.
Table 1: Comparison of GGA with the ERM baseline. The top out-of-domain accuracies on five domain generalization benchmarks averaged over three trials, are presented.
Algorithm PACS VLCS OfficeHome TerraInc DomainNet Avg.
ERM 85.5±$\pm 0.2$ 77.3±$\pm 0.4$ 66.5±$\pm 0.3$ 46.1±$\pm 1.8$ 43.8±$\pm 0.3$ 63.9
GGA (ours) 87.3±$\pm 0.4$ 79.9±$\pm 0.4$ 68.5±$\pm 0.2$ 50.6±$\pm 0.1$ 45.2±$\pm 0.2$ 66.3
Table 2: Comparison with state-of-the-art domain generalization methods. Out-of-domain accuracies on five domain generalization benchmarks are shown. Top performing methods are highlighted in bold while second-best are underlined. The results marked by $\dagger$,$\ddagger$ , ‡ are copied from Gulrajani and Lopez-Paz [31] and Wang et al. [81], respectively. For fair comparison, the training of each algorithm combined with GGA, were run on the respective codebases. Average accuracies and standard errors are calculated from three trials for the combination of GGA with past algorithms and from 5 trials for GGA. In green and red, we highlight the performance boost and decrease of applying GGA on top of each algorithm respectively, averaged over three trials. Due to computational resources, for DomainNet we do not combine GGA with previous methods.
Algorithm PACS VLCS OfficeHome TerraInc Avg. DomainNet Total
Mixstyle [95] 85.2±$\pm 0.3$ ($+0.3$) 77.9±$\pm 0.5$ ($+0.6$) 60.4±$\pm 0.3$ ($+0.5$) 44.0±$\pm 0.7$ ($+1.1$) 66.9 ($+0.6$) 34.0±$\pm 0.1$ 60.3
GroupDRO [66] 84.4±$\pm 0.8$ ($+1.4$) 76.7±$\pm 0.6$ ($+0.6$) 66.0±$\pm 0.7$ ($+2.2$) 43.2±$\pm 1.1$ ($+1.5$) 67.6 ($+1.4$) 33.3±$\pm 0.2$ 60.7
MMD [49] 84.7±$\pm 0.5$ ($+0.8$) 77.5±$\pm 0.9$ ($+1.3$) 66.3±$\pm 0.1$ ($+1.6$) 42.2±$\pm 1.6$ ($+6.3$) 67.7 ($+2.5$) 23.4±$\pm 9.5$ 58.8
AND-mask [69] 84.4±$\pm 0.9$ ($+0.1$) 78.1±$\pm 0.9$ ($+0.3$) 65.6±$\pm 0.4$ ($+1.2$) 44.6±$\pm 0.3$ ($-0.4$) 68.2 ($+0.3$) 37.2±$\pm 0.6$ 62.0
ARM [91] 85.1±$\pm 0.4$ ($+0.5$) 77.6±$\pm 0.3$ ($+0.9$) 64.8±$\pm 0.3$ ($+2.0$) 45.5±$\pm 0.3$ ($+0.8$) 68.3 ($+1.1$) 35.5±$\pm 0.2$ 61.7
IRM [1] 83.5±$\pm 0.8$ ($-1.5$) 78.5±$\pm 0.5$ ($-0.9$) 64.3±$\pm 2.2$ ($-2.1$) 47.6±$\pm 0.8$ ($-3.9$) 68.5 ($-2.1$) 33.9±$\pm 2.8$ 61.6
MTL [10] 84.6±$\pm 0.5$ ($+0.9$) 77.2±$\pm 0.4$ ($+2.0$) 66.4±$\pm 0.5$ ($+0.4$) 45.6±$\pm 1.2$ ($+0.6$) 68.5 ($+0.9$) 40.6±$\pm 0.1$ 62.9
VREx [42] 84.9±$\pm 0.6$ ($+0.6$) 78.3±$\pm 0.2$ ($+0.1$) 66.4±$\pm 0.6$ ($+1.3$) 46.4±$\pm 0.6$ ($+2.0$) 69.0 ($+1.0$) 33.6±$\pm 2.9$ 61.9
MLDG [48] 84.9±$\pm 1.0$ ($+0.7$) 77.2±$\pm 0.4$ ($+1.3$) 66.8±$\pm 0.6$ ($+1.2$) 47.7±$\pm 0.2$ ($+1.1$) 69.2 ($+1.2$) 41.2±$\pm 0.1$ 63.6
Mixup [84] 84.6±$\pm 0.6$ ($+1.2$) 77.4±$\pm 0.6$ ($+1.8$) 68.1±$\pm 0.3$ ($+1.0$) 47.9±$\pm 0.8$ ($+2.0$) 69.5 ($+1.5$) 39.2±$\pm 0.1$ 63.4
SagNet [56] 86.3±$\pm 0.2$ ($-1.0$) 77.8±$\pm 0.5$ ($+0.9$) 68.1±$\pm 0.1$ ($+0.3$) 48.6±$\pm 1.0$ ($+0.7$) 70.2 ($+0.4$) 40.3±$\pm 0.1$ 64.2
CORAL [72] 86.2±$\pm 0.3$ ($+0.7$) 78.8±$\pm 0.6$ ($-0.4$) 68.7±$\pm 0.3$ ($+0.2$) 47.6±$\pm 1.0$ ($+0.3$) 70.3 ($+0.2$) 41.5±$\pm 0.1$ 64.5
RSC [34] 85.2±$\pm 0.9$ ($+0.1$) 77.1±$\pm 0.5$ ($+0.2$) 65.5±$\pm 0.9$ ($+0.0$) 46.6±$\pm 1.0$ ($+0.2$) 68.6 ($+0.1$) 38.9±$\pm 0.5$ 62.7
Fish [71] 85.5±$\pm 0.3$ ($+0.1$) 77.8±$\pm 0.3$ ($+0.9$) 68.6±$\pm 0.4$ ($-0.6$) 45.1±$\pm 1.3$ ($+3.8$) 69.3 ($+1.0$) 42.7±$\pm 0.2$ 63.9
SAM [28] 85.8±$\pm 0.2$ ($+0.6$) 79.4±$\pm 0.1$ ($+0.7$) 69.6±$\pm 0.1$ ($+0.4$) 43.3±$\pm 0.7$ ($+2.6$) 69.5 ($+1.1$) 44.3±$\pm 0.0$ 64.5
GSAM [98] 85.9±$\pm 0.1$ ($+0.4$) 79.1±$\pm 0.2$ ($+1.0$) 69.3±$\pm 0.0$ ($+0.3$) 47.0±$\pm 0.8$ ($+0.6$) 70.3 ($+0.2$) 44.6±$\pm 0.2$ 65.1
SAGM [81] 86.6±$\pm 0.2$ ($+0.2$) 80.0±$\pm 0.3$ ($-0.3$) 70.1±$\pm 0.2$ ($-0.6$) 48.8±$\pm 0.9$ ($-0.1$) 71.4 ($-0.2$) 45.0±$\pm 0.2$ 66.1
GGA (ours) 87.3±$\pm 0.4$ 79.9±$\pm 0.4$ 68.5±$\pm 0.2$ 50.6±$\pm 0.1$ 71.7 45.2±$\pm 0.2$ 66.3
4 Experiments
4.1 Experimental setup and implementation details
In our experiments, we follow the protocol of DomainBed [31] and exhaustively evaluate our algorithm against state-of-the-art algorithms on five DG benchmarks, using the same dataset splits and model selection. For the hyperparameter search space, we follow [15] in order to avoid the high computational burden of DomainBed. The datasets included in the benchmarks are, PACS [47] (9,991 images, 7 classes, and 4 domains), VLCS [26] (10,729 images, 5 classes, and 4 domains), OfficeHome [75] (15,588 images, 65 classes, and 4 domains), TerraIncognita [7] (24,788 images, 10 classes, and 4 domains) and DomainNet [59] (586,575 images, 345 classes, and 6 domains).
In all experiments, the leave-one-domain-out cross-validation protocol is followed. Specifically, in each run a single domain is left out as the target (test) domain, while the rest of the domains are used for training. The final performance of each algorithm is calculated by averaging the top-1 accuracy on the target domain, with different train-validation splits. For training, we utilize a ResNet-50 [32] pretrained on ImageNet [64] for the backbone feature extractor and ADAM for the optimizer. Regarding GGA, during training we let each algorithm run for several training iterations depending on the dataset and then begin the parameter space search, as described in Section 3. For the neighborhood size in the search process, we set $\rho$ to 0.00050.00050.00050.0005333The selection of $\rho$ was based on previous algorithms that implement weight perturbations in ResNet-50 networks [28, 81, 45]. The sensitivity analysis presented in subsection 4.4 also reveals that $\rho=0.0005$ yields the best performance. and search for a total of $A=250$ steps before moving to the next mini-batch of 48 samples from each domain, in each dataset. In all experiments, we perform search iterations for $100$ different mini-batches during early training stages. The rest of the hyperparameters, such as learning rate, weight decay and dropout rate, are tuned according to [15] and are presented in Table 3. To account for the variability introduced in the random search of GGA, we repeat the experiments with 5 different seeds for each dataset. All models were trained on a cluster containing $4\times 40$GB NVIDIA A100100100100 GPU cards, split into $8$ $20$GB virtual MIG devices and $1\times 24$GB NVIDIA RTX A5000500050005000 GPU card.
Hyperparameter PACS VLCS OH TI DN
Learning rate 3e-5 1e-5 1e-5 1e-5 3e-5
Dropout 0.5 0.5 0.5 0.5 0.5
Weight decay 1e-4 1e-4 1e-4 1e-4 1e-4
Training Steps 5000 5000 5000 5000 15000
$\rho$ 5e-5 5e-5 5e-5 5e-5 5e-5
GGA Start-End 100-150, 1500-1550 100-200 100-200 500-600 100-200
Table 3: Hyperparameters for DG experiments. OH, TI and DN stand for OfficeHome, TerraIncognita and DomainNet respectively. $\rho$ is the parameter space search used during the annealing steps in GGA, while the last row indicates the training iterations during which GGA occurs.
4.2 Comparative Evaluation
The average OOD performances of the baseline vanilla ERM [74] and state-of-the-art DG methods on a total of 5 DG benchmarks, are reported in Tables 1 and 2 respectively. The results for each separate domain in each benchmark are reported in the Appendix. In the experiments, we apply GGA on-top of the rest of the DG algorithms and show that its properties generalize to other methods as well, boosting their overall performance. It should also be noted that GGA can be adapted to training scenarios where distributions shifts are present in training data.
Initially, to validate whether the application of GGA in the early stages of model training leads to models with increased generalizabilty, we compare the results with the vanilla ERM baseline. As presented in Table 1, GGA is able to boost the performance of the baseline model by an average of 2.4% on all benchmarks, demonstrating the efficacy of the proposed algorithm. For further evaluation, we also compare GGA with state-of-the-art DG methods [49, 95, 66, 1, 91, 42, 69, 30, 34, 10, 84, 48, 71, 56, 72, 28, 98, 81, 74], before applying GGA to them as well. We note that we only include previous works that have implemented a ResNet-50 as the backbone encoder and do not use additional components or ensembles. In Table 2 we differentiate between the methods that operate on model gradients [34, 71, 28, 98, 81] and the rest of the algorithms. Even without its application to other methods, GGA is able to surpass most of the previously proposed algorithms, while also setting the state-of-the-art on PACS ($+0.7\%$) and on the challenging TerraIncognita ($+1.8$) dataset, while remaining competitive in the other two. What’s more important is that the application of GGA in conjunction with the rest of the DG methods, proves beneficial and ultimately boosts their overall performance in almost each case by an average of around $1\%$ and in some cases up to even $6.3\%$. With regards to IRM, which seems to be significantly impacted negatively by the application of GGA, its learning objective emphasizes on simultaneously minimizing the training loss of each source domain and not necessarily on the pairwise agreement of gradients among domains. It is therefore not able to converge to a good solution after the initial model’s weights have been perturbed during training.
From the experimental results, it is evident that the application of GGA and its search for parameter values where gradients align between domains is beneficial to model training. By introducing the proposed annealing step before the final stages of training, the majority of the models exceed their previous performance and exhibit improved generalization capabilities.
Figure 3: Impact of GGA on gradient alignment during model training on the VLCS dataset. As illustrated, in the case of vanilla ERM, even though the training loss is minimized, the average cosine similarity among domain gradients remains low. In the case of GGA however, after the algorithm searches for points in the parameter space with increased gradient similarity, the gradients continue to agree during training, while the total training loss is also minimized.
4.3 Evaluating the impact of GGA on gradient disagreement
As discussed in Section 3, when a training dataset is composed as a mixture of multiple domains, conflicting gradients between mini-batches drawn from each domain lead to models that do not infer based on domain-invariant features and which generalize to previously unseen data samples, but are hindered by domain-specific, spurious correlations. This is evident in the case of vanilla models trained via ERM where the average gradient similarity among domains continues to remain low upon reaching a local minima. Our hypothesis is that this behavior can be avoided by searching for a parameter set of common agreement between domains before optimizing via gradient descent.
To demonstrate the operation of the proposed algorithm in practice against ERM, we calculate the average gradient cosine similarity between mini-batches from source domains during training for the VLCS dataset, along with the training loss in each iteration. As a result, each sub-figure in Figure 3 illustrates the progression of the training gradient alignment between domains, against the total training loss.
As expected, in the very initial iterations the gradients of the pretrained model parameters point towards a common direction. However, in the case of ERM as training progresses and the loss is minimized, the domain gradients begin to disagree leading the model to converge to undesirable minima that do not generalize across domains. On the other hand, when GGA is applied the model searches for parameters such that gradients are aligned before continuing training. This is illustrated by the spike in gradient similarity, during iterations $100$ up to $200$. After GGA concludes, we observe that the model continues training by descending into minima where gradients agree among domains.
4.4 Sensitivity Analysis
Figure 4: Sensitivity analysis of the parameter space search magnitude $\rho$ and the training stage application of GGA. The analysis on PACS reveals smaller weight perturbations and gradient annealing during early training iterations lead to increased model performance.
The two core parameters of GGA, are the size of the parameter search $\rho$ and the moment of our methods implementation during training, i.e., during the early, mid or late training stages. To justify their selection, we conduct a sensitivity analysis (Fig. 4) by varying one of the above parameters while fixing the other at its optimal value.
Regarding the magnitude of weight perturbations during the application of GGA, we found that the optimal value was $\rho=5e-5$. As illustrated in Figure 4, a larger magnitude of perturbation led to decreased model performance. Intuitively, the application of larger noise to the model parameters leads to sets that are not close to the solution, making it increasingly difficult for the model to converge. On the other hand, smaller perturbations seem to have little to no effect on training, as the search is limited to spaces near the current parameters, which is why the model performance falls back close to that of ERM. With regards to the stage of training during which GGA will be applied, we found that the models yielded better performance when the search was initialized in earlier stages.We hypothesize that applying perturbations near the end of training displaces the model from a local optimum, requiring additional iterations to converge.
5 Conclusions
In this work, we investigated gradient conflicts in the context of domain generalization, leading to the development of GGA; a simple yet effective algorithm that identifies points in the parameter space where domain gradients align early in training before continuing optimization, ultimately enhancing model generalization. Through a comprehensive comparative analysis we have demonstrated the efficacy of GGA, which is able to achieve superior generalization capabilities in standard DG benchmarks and outperform SoTA baselines. What’s more, as GGA is both model and method agnostic it can also be utilized by other algorithms. Interestingly, its combination with previous methods proves beneficial, as their majority yields improved performance over all benchmarks. Finally, we validate the impact of our algorithm on gradient alignment through additional experiments, showing that, unlike the baseline, domain gradients align during training.
However, our method does not come without some key limitations. First of all, the annealing steps and gradient similarity computation add computational overhead. Furthermore, GGA relies on domain labels, making it inapplicable to single-source DG, where domain labels are not available during training. Its effectiveness is also batch-size dependent, as larger batches provide further information regarding gradient alignment and can lead to improved results. Future work will further explore gradient alignment in multi-domain settings and refine GGA to make annealing steps adaptive rather than random.
Acknowledgement. This work was supported by the European Union’s Horizon 2020 research and innovation programme under Grant Agreement No. 965231, project REBECCA (REsearch on BrEast Cancer induced chronic conditions supported by Causal Analysis of multi-source data).
Supplementary Material
The following materials are provided in this supplementary file:
  • An extended literature review discussion, helpful for navigating the Domain Generalization literature under the scope of computer vision.
  • A computational analysis regarding the application of GGA.
  • Detailed results for each dataset domain and algorithm, presented in Table 2 of the main text, along with extra experiments on the ColoredMNIST and RotatedMNIST datasets.
F Extended Literature Review
Domain Generalization (DG) [79] is arguably one of the most difficult and fundamental problems of Machine Learning (ML) today. Unsurprisingly, a vast number of researchers have poured effort into advancing the field, where findings have been applied to various areas, such as Natural Language Processing [35], Reinforcement Learning [48], Healthcare and Medicine [23, 4], Time-Series forecasting [25], Fault Diagnosis [93] and, of course, Computer Vision [79]. Even though not covering the entire field of DG, this section aims to present a taxonomy of the general DG methodologies developed in CV, for producing robust models that can generalize to previously unseen data, and attempts to assist potential readers navigate the past literature, while also categorizing our proposed method among its predecessors. Domain Generalization methods can be categorized into three major groups, depending on their operation during the process of model training, namely: (a) Data Manipulation, (b) Representation Learning and, (c) Learning Algorithm. Furthermore, as mentioned in the main text, DG algorithms can either leverage domain labels during training (multi-source), or completely disregard the knowledge of existing domain shifts in their training data and handle them as a single distribution (single-source).
Data Manipulation. As its name suggests, methods included in this group focus on either perturbing existing samples (data augmentation) or creating novel ones (data generation), in order to regularize the training of machine learning models, avoid overfitting and improve their generalizability. The basic idea in data manipulation methodologies is to simulate domain shift by creating diverse data samples, which can in turn mimic the entirety of distributions present in the input space. Regarding data augmentation, most popular techniques include traditional image transformations, such as random flip, rotation and color distortion. Even though these augmentations can be randomly applied during training, without needing domain labels, it has been shown that their selection significantly affects model performance. For example, the authors of [76] define novel augmentation rules that push the perturbed images to diverge as much as possible from the original ones. Additionally, image augmentations prove effective towards overcoming domain shifts in medical image classification [58, 90], where transformations can replicate shifts caused by the use of different devices. On the other hand, multiple data augmentation methods were also inspired by adversarial attacks and use adversarial gradients to distort the input images [77, 62], or use randomly initialized convolutional networks for transforming samples [19]. These techniques act as regularizers during model training, allowing them to learn generalizable image representations. The generation of novel data domains is also a well researched area in the data manipulation group. In addition to using domain gradients for synthesizing novel domains [70], several methods took advantage of style transfer [33] and either map the styles of images to that existing source domains [11] or create novel styles [87]. On a similar note, mixing the styles of training images by conventional methods [84, 95] or with the generative models [80] also proves beneficial.
Representation Learning. This group of methods is arguably the most prominent in DG and has been the central focus of ML [9]. Following the formulation in the main text, given a labeling function $h$ that maps input observations $\bm{x}$ to their labels $y$, we can decompose it into $h=f\circ g$, where $g$ is a parametric function that learns representations of $x$ and $f$ is the classifier function. The goal of representation learning can be summarized as follows:
$\min_{f,g}\mathbb{E}_{x,y}\ell(f(g(\bm{x};\bm{\theta})),y)+\lambda\ell_{\text{reg}}$ (6)
where $\ell$ the loss function to be minimized and $\ell_{\text{reg}}$ a regularizer. Methods included in this group, focus on learning a robust and generalizable representation learning function $g$. The algorithms included in this group can be further categorized into three sub-groups. Feature disentanglement [88] methods intend to extract disentangled feature vectors from samples, where each dimension can be linked to a subset of data generating factors. The main idea is to produce a model that extracts a representation that can be further decomposed into domain-specific, domain-invariant, and class-specific features. To that end, the authors of [61] present CSD, which jointly learns a domain-invariant and domain-specific component in the final embedding and enables the extraction of disentangled representations, whereas the authors of [16] propose learning domain specific masks during training to improve model robustness. Generative models have also been proposed in the disentangled representation learning literature for DG, with variational autoencoders (VAEs) and GANs [17] being utilized for learning distinct latent subspaces for class- and domain-specific features [36]. Another promising category of methods aiming to produce disentangled representations is that of Causality-Inspired algorithms. In causal representation learning, a domain shift can be thought of as an intervention, subsequentially leading the development of models that aim to uncover the true causal data generating factors. Naturally, the prediction of a model should not be affected by interventions on spuriously correlated but irrelevant features, such as the background, color or style of the image. Under this causal consideration, the authors of [52] propose a structural causal model in order to model within-class variations and leverage the fact that inputs across domains should have the same representation, given that they derive from the same object. Similar to disentangled representations, there have been proposed methods in the literature that focus on completely disregarding domain-relevant from the final feature vectors, deriving solely domain-invariant representations. Based on the initial findings of [8], numerous works have presented algorithms that aim to minimize the representation differences across multiple source domains within a specific feature space, ensuring they become domain invariant, ultimately enabling the trained model to effectively generalize to previously unseen domains. In one of the most notable previous works in this category, Arjovsky et al. [1] enforce the optimal classifier on top of the representation space to be the same across domains and simultaneously minimize the loss across distributions. The above idea of Invariant Risk Minimization (IRM) has been extended to several other works. For example, the authors of [43] propose minimizing the variance of source-domain risks, by minimizing their extrapolated risk, while the authors of [91] propose adapting to domain shift and producing invariant representations. Finally, an alternative route towards learning generalized representations is via regularization strategies. The most representative group of methods in this category is Gradient-Based operations, which utilize gradient information during model training. In [34], the authors propose learning robust representations by discarding the most dominant gradients in each training iteration under the assumption that they are correlated with domain-specific features present in the source data. Another popular strategy is to seek for flat minima [28, 15] in the loss landscape of neural networks during training, assuming that models that converge to flat minima exhibit increased generalization capabilities [98, 81]. What’s more, Shi et al. [71] hypothesized that gradients among domains should match and proposed an approximation of a loss inducing the maximization of the gradient inner product during training. Our method (GGA) can be categorized in this group of gradient operations, as it considers the similarity of domain gradients in the early iterations of model training and seeks for sets in the parameter space with increased gradient alignment, before continuing the optimization procedure.
Learning Algorithm. In addition to manipulating the input space or feature extractor, DG methods were also researched under the scope of alternative ML learning paradigms, such as ensemble, meta, domain-adversarial, self-supervised and reinforcement learning. In this section we present the most exemplary works in each category. Ensemble-Learning in DG initially combined several copies of the same network, each of which is trained on a specific domain [94, 22]. Alternatively, instead of using several networks, [85] proposed sharing shallow layers among CNNs. During inference, the final prediction is produced by either simple [94] or weighted averaging [82]. In Meta-Learning for DG, Li et al. [48] propose MLDG and split the source domains into meta-train and meta-test splits to mimic the effects of domain shift during training. Similarly, [3] proposes learning a meta regularizer for the classifier, while MAML [27] was proposed for improved parameter initialization. Another approach is that of Adversarial Learning (AL). In the context of DG, the aim of adversarial learning is to train a classifier to distinguish between source domains [54] and ultimately learn domain-agnostic features from the samples that can be generalized to unseen data [50]. Other learning paradigms such as Self-Supervised learning have also been explored in DG, which leverages unlabeled data samples to derive generalized representations. Notably, the authors of [14] introduce a self-supervised jigsaw-solving puzzle task to push the model to learn robust representations. Furthermore, contrastive learning has also been shown to improve model performance. Specifically, SelfReg [39] utilizes self-supervised contrastive losses to bring latent representations of same-class samples closer. Similarly, the authors of [6] introduce a contrastive loss for representations extracted from intermediate layers of the network. Finally, Reinforcement learning has also been applied in the context of DG. Indicatively, previous works have explored randomizing the environments of an RL agent for transferring them to real-world scenarios [73, 46], whereas [44] researches the combination of RL with contrastive learning.
G Computational Analyis
G.1 Experiment Infrastructure
Each and every experiment is conducted on a cluster containing $4\times 40$GB NVIDIA A100100100100 GPU cards, split into $8$ $20$GB virtual MIG devices and $1\times 24$GB NVIDIA RTX A5000500050005000 GPU card, via a SLURM workload manager.
G.2 Complexity Analysis
Each GGA training iteration includes computing model gradients $S\cdot n_{a}$ times for each training step, where $S$ is the number of source domains and $n_{a}$ is the number of search steps. These GGA training iterations only take place in the early stages of training and for a small percentage of the total training iterations (2% in our experiments). The rest of the iterations are vanilla ERM. Furthermore, inference is not affected by the application of GGA during training.
H Full Experimental Results
In this section, we show detailed results of Table 2 in the main text. The results marked by $\dagger$,$\ddagger$ , ‡ are copied from Gulrajani and Lopez-Paz [31] and Wang et al. [81], respectively. Standard errors for the baseline methods are reported from three trials, if available from past literature. In green and red, we highlight the performance boost and decrease of applying GGA on top of each algorithm respectively, averaged over three trials. In addition, we also present detailed results for the DomainNet benchmark, without however including results for the combination of GGA with the baseline algorithms, due to computational restrictions. We also include experiments for the ColoredMNIST and RotatedMNIST datasets, where we reproduced the results for all baselines and report the average results over 5 runs. The below tables are better read in color.
When applying GGA to existing methods, the only difference regarding the baseline algorithm training is that “Algorithm 1” (i.e. GGA) is applied instead of the method’s update rules for the duration of the annealing process (training steps $A_{s}$ to $A_{e}$). The total epochs and method hyperparameters remain the same throughout training.